What is aceptable drawing accuracy standard?
The answer depends on the purpose. For example, the accuracy required from an engineering drawing is much greater than a child's drawing of his house.
How do the principels of probability explain the results of Mendel's genetic corsses?
this is not fair you say ask anything and when i do you only can answer half of my questions
What is the probability of pulling a king of hearts out of a deck?
The probability of pulling a king of hearts in one random selection from a well shuffled normal deck of cards is 1/52.
What proportion of a normal distribution is located between z -0.50 and z 0.50?
Prob(-0.5 < z < 0.5) = 0.3830
How does the mean affect the graph of a normal distribution?
It determines the location of the graph: left or right - but not its shape.
Number of elements int the event space of rolling a sum of 7 with two dice?
When you roll two dice, a 7 may occur as (1,6),(2,5),(3,4),(4,3),(5,2),(6,1)
There are 6 elements in the event space.
The probability that five cards chosen at random from a standard deck will all be the same suit is
12/51 * 11/50 * 10/49 * 9/48 = about 0.2%.
When does normal distribution occur?
The normal distribution occurs when a number of random variables, with independent distributions, are added together. No matter what the underlying probability distribution of the individual variables, their sum tends to the normal as their number increases.
Many everyday measures are composed of the sums of small components and so they follow the normal distribution.
What is the theoretical probability of a six sided die when it is rolled 24 times?
well its all 50% no matter what
What is the probability of rolling a 4 or an odd number on a die?
The probability of rolling a 4 is 1/6. (There are 6 numbers one of which is "4")
The probability of rolling an odd number is 3/6 = 1/2 (There are 6 numbers 3 of which are odd)
Thus, in one roll of the die, the odds of rolling a 4 oran odd number = 1/6 + 3/6 = 4/6 = 2/3.
What is probability of drawing a diamond and a club from a deck of cards?
The answer depends on how many cards are drawn. It is 0 if only one card is drawn and 1 if 40 cards are drawn without replacement.
What is the probability of getting three heads in tossing a coin three times?
It is (1/2)3 = 1/8 or 0.125
What percentage do you have of winning top prize on a scratch card?
There are loads of scratch cards across the world and the winning probabilities are all different.
Is the standard deviation best thought of as the distance from the mean?
No. A small standard deviation with a large mean will yield points further from the mean than a large standard deviation of a small mean. Standard deviation is best thought of as spread or dispersion.
What would be the theoretical probability of rolling a die 50 times and rolling a 2?
The probability of rolling at least one 2 in fifty rolls of a standard die is 1 - (5/6) 50, or about 0.99989012.
This calculation starts by looking at the probability of not rolling a 2, which is 5/6. To repeat that 50 times in a row, you simply raise that to the 50th power, getting 0.000109885. Then you subtract the result from 1 to get the probability of not succeeding in not rolling a 2 in fifty tries.
Expressed in normal "odds" notation, this is about (100000 - 11) in 100000, or about 99989 in 100000.
What is the probability of choosing a marble that is not green in problem 2?
Since you haven't seen fit to tell us problem 2, we'd only be guessing.
What are the possible outcomes of tossing a quarter and a dime while rolling a number cube?
it would be 2*2*6 which would equal 24
How do theoretical probability and experimental probability relate?
Take for example, flipping a coin. Theoretically, if I flip it, there is a 50% chance that I flip a head and a a 50% chance that I flip a tail. That would lead us to believe that out of 100 flips, there should theoretically be 50 heads and 50 tails. But if you actually try this out, this may not be the case. What you actually get, say 46 heads and 54 tails, is the experimental probability. Thus, experimental probability differs from theoretical probability by the actual results. Where theoretical probability cannot change, experimental probability can.
The probability of drawing a heart from a fair deck is 1 in 4. If the card is replaced then the probability is again 1 in 4.
The probability of drawing a card other than a heart is 3 in 4. Once again if the card is replaced then the probability remains 3 in 4
What is the probability of choosing a red spade?
The probability of drawing a red spade is zero. There are no red spades in a standard deck.